5-18. The following data are obtained from the infrared spectrum of 1271CI. Using the method of Problem 5-17, determine the values ofae and, from these data. Transition Frequency/cm1 381.20 759.60 1135.00 1507.40 1877.00 0?4
Answer #1

Alright starting with 5.17, the energy required for the molecule jump any vibrational level from v = 0, to v = 1,2,3,...

\tilde{\omega}_{obs} = G(v) - G(0) = \tilde{\omega}_ev - \tilde{\chi}_e\tilde{\omega}_ev(v+1)

Dividing the equation with v on both sides,

\frac{\tilde{\omega}_{obs}}{v} = \tilde{\omega}_e - \tilde{\chi}_e\tilde{\omega}_e(v+1)

Which can be compared to the equation of a line, y = mx+c with m slope and c intercept.

This gives us the value of what we need very conveniently,

\tilde{\chi}_e\tilde{\omega}_e = -m
\tilde{\omega}_e = c

It is apparent that since it is anharmonic oscillator, as the wave-number rises, the energy gaps between two levels starts decreasing up to a point where a continuum is obtained.

(5.18) Let us first write the data we need,

Data we need
v+1 \tilde{\omega}_{obs}/v
2 381.20
3 379.80
4 378.33
5 376.75
6 375.40

The graph I've plotted is a crude one on some online plotting website. A better job can be there, but here will work just fine.

Click to enter Plot title 384 382 380 378 376 374 372 V+1Calculating slope should be easy since the line quite linear,

m = -\frac{381.20-375.4}{6-2} = -1.45

\therefore -m = \tilde{\chi}_e\tilde{\omega}_e = 1.45 \ cm^{-1}

The intercept can be extrapolated as,

c \sim 384\ cm^{-1} = \tilde{\omega}_e